A 74-g ice cube at 0°C is placed in 841 g of water at 30°C. What is the final temperature of the mixture?
Answer #1

Energy absorbed by ice = mass * Heat of fusion = 74 * 334 = 24,716 Joules

Energy released by water = Mass * Specific heat * delta T = 841 * 4.18 * delta T

841 * 4.18 * delta T = 74 * 334

delta T = 7.02˚

Final temperature of water = 30 - 7.02 = 22.98

The final temperature of the water

74 * 4.186 * (Tf – 0) = 841 * 4.186 * (22.98 – Tf)

74 * Tf = 841 * (22.98 – Tf)

915 * Tf = 19326

Tf = 21.12